OpenStudy (anonymous):

Write the equation y - 2 = 4(x + 5) in standard form. A. 4x + y = 7 B. 4x - y = 3 C. 4x - y = -22

OpenStudy (anonymous):

please

OpenStudy (anonymous):

y-2=4(x+5), y-2=4x+20, y-2+2=4x+20+2, y=4x+22, -4x+y=22, -1(-4x+y=22), 4x-y=-22. The answer is C. Does this make sense?

OpenStudy (anonymous):

sort of.

OpenStudy (anonymous):

how did it become negative?

OpenStudy (anonymous):

I had to multiply the whole problem by -1 because the variable in the problem (-4x) was not supposed to be negative. -1(-4x+y=22), see? and then your answer is 4x-y=-22.

OpenStudy (anonymous):

ohh i see.thank you.

OpenStudy (anonymous):

Not a problem ^_^

OpenStudy (anonymous):

P.S. Oh you may want to post this in the math section instead of the English next time.